Suppose you invest $15000 at 6%. How much additional money should you invest at 4% to produce a yield of 5.5% on the entire amount invested?

Answer by richwmiller(17219)

one investment all 3 rates
0.06*15000+0.04(x)=0.055(15000+x)
Multiply out.
900+0.04(x)=825.0+0.055x
combine like terms
-0.015x=-75.0
isolate x
x=-75.0/-0.015
Divide
x=5000.0
check
0.06*15000+0.04(x)=0.055(15000+x)
900+0.04*5000=0.055(15000+5000)
900+200=0.055(20000)
1100=1100.0
ok